OpenStudy (anonymous):

Which of the following choices is the explicit formula for the following geometric sequence? -0.1, 0.04, -0.016, 0.0064, -0.00256, ... A. an = 0.1(0.4)n B. an = 0.1(-0.4)n-1 C. an = 0.4(-0.4)n-1 D. an = -0.1(-0.4)n-1 E. an = -0.1(0.4)n-1

OpenStudy (anonymous):

@magepker728 could you help me out one more time?

OpenStudy (anonymous):

or @ShadowLegendX

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

do you know the original formula ??

Nnesha (nnesha):

well its \[\huge\rm a _{n}=a _{1} \times r ^{n-1}\]

Nnesha (nnesha):

where a1 us first term and r is ratio

OpenStudy (anonymous):

r is also in parantheses as well as n-1

Nnesha (nnesha):

n-1 is exponent of r

Nnesha (nnesha):

first tell me what is first term ??

OpenStudy (anonymous):

an=-0.1?

Nnesha (nnesha):

yes not an a1 = -0.1

Nnesha (nnesha):

did you notice first term is negative and second term is positive so what your ratios suppose to be(all answer choices have same ratio but different sign ) so see which one is right is it -.4 or positive .4 ??

OpenStudy (anonymous):

it should be a positive 0.04

Nnesha (nnesha):

yes right so replace r by 0.4 and a1 by -0.1 in the formula that i gave you

OpenStudy (anonymous):

an=0.1(0.04)^(n-1) right?

Nnesha (nnesha):

nope your first term is negative

OpenStudy (anonymous):

an=-0.1(0.04)^(.-1) how bout now?

Nnesha (nnesha):

anh

Nnesha (nnesha):

nope

Nnesha (nnesha):

exponent is n-1

OpenStudy (anonymous):

(n-1) i think i hit the wrong button on that one

Nnesha (nnesha):

yes then that's right

OpenStudy (anonymous):

so it would be D

Nnesha (nnesha):

yes right

Nnesha (nnesha):

wait

Nnesha (nnesha):

nope

OpenStudy (anonymous):

actually not. its gotta be both negative

Nnesha (nnesha):

yeah that's right ^

Nnesha (nnesha):

btw yes it is D i thought you just have 4 choices and D is E but there is 5 answer chocies

Nnesha (nnesha):

@Tallan so yeas D is right answer bec ratio is negative .4

Nnesha (nnesha):

@Tallan did you get this one

OpenStudy (anonymous):

i got it yea... Could you help me with the other problem i posted?

Nnesha (nnesha):

okay :)